# USV-AUV Collaboration Framework for Underwater Tasks under Extreme Sea Conditions

### v2

@@ Abstract
Autonomous underwater vehicles (AUVs) are valuable for ocean exploration due to their flexibility and ability to carry communication and detection units. Nevertheless, AUVs alone often face challenges in harsh and extreme sea conditions. This study introduces a unmanned surface vehicle (USV)–AUV collaboration framework, which includes high-precision multi-AUV positioning using USV path planning via Fisher information matrix optimization and reinforcement learning for multi-AUV cooperative tasks. Applied to a multi-AUV underwater data collection task scenario, extensive simulations validate the framework’s feasibility and superior performance, highlighting exceptional coordination and robustness under extreme sea conditions. To accelerate relevant research in this field, we have made the simulation code available as open-source1.
@@ I Introduction
Current positioning techniques for AUVs mainly consist of global positioning systems, inertial navigation systems, and ultrasonic positioning. However, these methods are vulnerable to problems like water refraction and scattering, signal attenuation, transmission delays [3, 8], and error accumulation [5]. In addition, traditional control strategies for multi-AUV systems depend heavily on mathematical models or model-based algorithms [9]. Model-based approaches face difficulties when controller parameters shift in dynamic environments [10], and both algorithms often struggle to predict the future maneuvers and behaviors of non-cooperative targets, thereby limiting their scalability and adaptability [9].
@@ IV Conclusion
In this paper, we propose a USV–AUV collaboration framework. The two parts of our framework, including high-precision multi-AUV location using USV path planning by FIM optimization and RL training for multi-AUV cooperative tasks, collaboratively enhance the performance of multi-AUV underwater tasks in extreme sea conditions. Experiment results from underwater data collection task verify the superior feasibility and performance of the proposed framework, which effectively embodies high coordination between USV and AUV while showcasing excellent robustness in extreme sea conditions. To accelerate relevant research in this field, we have made the simulation codes available as open-source at https://github.com/360ZMEM/USV-AUV-colab.

### v3

@@ Abstract
Autonomous underwater vehicles (AUVs) are valuable for ocean exploration due to their flexibility and ability to carry communication and detection units. Nevertheless, AUVs alone often face challenges in harsh and extreme sea conditions. This study introduces a unmanned surface vehicle (USV)–AUV collaboration framework, which includes high-precision multi-AUV positioning using USV path planning via Fisher information matrix optimization and reinforcement learning for multi-AUV cooperative tasks. Applied to a multi-AUV underwater data collection task scenario, extensive simulations validate the framework’s feasibility and superior performance, highlighting exceptional coordination and robustness under extreme sea conditions. To accelerate relevant research in this field, we have made simulation code (demo version) available as open-source1.
@@ I Introduction
Current positioning techniques for AUVs mainly consist of global positioning systems, inertial navigation systems, and ultrasonic positioning. However, these methods are vulnerable to problems like water refraction and scattering, signal attenuation, transmission delays [3, 8, 9], and error accumulation [5]. In addition, traditional control strategies for multi-AUV systems depend heavily on mathematical models or model-based algorithms [10]. Model-based approaches face difficulties when controller parameters shift in dynamic environments [12], and both algorithms often struggle to predict the future maneuvers and behaviors of non-cooperative targets, thereby limiting their scalability and adaptability [10, 11].
@@ IV Conclusion
In this paper, we propose a USV–AUV collaboration framework. The two parts of our framework, including high-precision multi-AUV location using USV path planning by FIM optimization and RL training for multi-AUV cooperative tasks, collaboratively enhance the performance of multi-AUV underwater tasks in extreme sea conditions. Experiment results from underwater data collection task verify the superior feasibility and performance of the proposed framework, which effectively embodies high coordination between USV and AUV while showcasing excellent robustness in extreme sea conditions. To accelerate relevant research in this field, we have made simulation codes (demo version) available as open-source at https://github.com/360ZMEM/USV-AUV-colab .
V Acknoledgments
Part of this work was done when Jingzehua Xu was studying at Tsinghua University, and in the MicroMasters Program in Statistics and Data Science at Massachusetts Institute of Technology (MIT), respectively. We would like to express our heartfelt gratitude to Yong Ren at Tsinghua University for his generous praise. Our sincere appreciation also goes to Prof. Shuai Zhang at the New Jersey Institute of Technology (NJIT) for his unwavering support. Furthermore, we genuinely appreciate the encouragement from Prof. Miao Liu and Prof. Songtao Lu at IBM Research, as we have gained numerous valuable insights from their previous work, and their positive feedback has greatly inspired us to move forward in multi-agent systems.